# Interpretability and Generalization Bounds for Learning Spatial Physics

### v1

@@ Abstract
While there are many applications of ML to scientific problems that look promising, visuals can be deceiving. For scientific applications, actual quantitative accuracy is crucial. This work applies the rigor of numerical analysis for differential equations to machine learning by specifically quantifying the accuracy of applying different ML techniques to the elementary 1D Poisson differential equation. Beyond the quantity and discretization of data, we identify that the function space of the data is critical to the generalization of the model. We prove generalization bounds and convergence rates under finite data discretizations and restricted training data subspaces by analyzing the training dynamics and deriving optimal parameters for both a white-box differential equation discovery method and a black-box linear model. The analytically derived generalization bounds are replicated empirically. Similar lack of generalization is empirically demonstrated for deep linear models, shallow neural networks, and physics-specific DeepONets and Neural Operators. We theoretically and empirically demonstrate that generalization to the true physical equation is not guaranteed in each explored case. Surprisingly, we find that different classes of models can exhibit opposing generalization behaviors. Based on our theoretical analysis, we also demonstrate a new mechanistic interpretability lens on scientific models whereby Green’s function representations can be extracted from the weights of black-box models. Our results inform a new cross-validation technique for measuring generalization in physical systems. We propose applying it to the Poisson equation as an evaluation benchmark of future methods.
@@ 1 Introduction
The development of robust and rigorous a priori estimates for many numerical methods is a major victory of the past century. These estimates ensure geometric flexibility and convergence guarantees, laying the groundwork for complex physics simulations across a wide range of applications (Ciarlet, 2002; LeVeque, 2007). The foundation of all these results is the humble 1D Poisson equation, studied by virtually all numerical analysts and serving as the theoretical bedrock for computational physics.
Data-driven machine learning has risen dramatically recently as a new way to model previously intractable systems (Raissi et al., 2019; Lu et al., 2019). While ML/SciML is already successfully deployed in disparate domains from autonomous robotics to climate simulations, its theoretical foundations are crucial. Scientific applications are unique in the need for numerical precision and accuracy: even small percentage errors in parameters can be impactful. Recent works have advanced the theoretical understanding learning time-dependent differential equations (Queiruga et al., 2020; Ott et al., 2021; Zhu et al., 2022; Krishnapriyan et al., 2023), but thus far no comprehensive theory exists for learning the spatial component. Just as a foundational textbook chapter on classical numerical analysis rigorously studies solving the Poisson equation, this paper rigorously studies learning the Poisson equation to build the same theoretical foundations for ML. We combine traditional theoretical and empirical analysis with modern lenses, such as mechanistic interpretability through direct parameter inspection.
Figure 1: Expectations when learning a black box linear model, ${\bm{u}}={\bm{W}}{\bm{f}}$. Despite some visual similarity to the Green’s function operator (top), naive approaches are not guaranteed to converge to the true operator. The issue lies in the sampling procedure of the training data. The naive approach of using polynomial analytical solutions (illustrated as cubic polynomials) can achieve machine precision MSE loss, but even with infinitely many examples, will not converge to the general solution (center). It is possible to alter the construction of the training data, i.e. by using random piecewise linear functions solved via the finite element method (FEM) (bottom) to recover the true operator. This further allows for the extraction of a familiar discrete differential equation operator (right).
The canonical 1D Poisson equation on the unit interval with homogeneous Dirichlet boundary conditions is given by
$-k\frac{d^{2}u}{dx^{2}}=f(x),\qquad u(0)=u(1)=0.$ (1)
where $u(x)$ the solution field, $f(x)$ the forcing function, and $k$ the material property, e.g. thermal conductivity. The solution to Eq. (1) can be obtained by Green’s function which maps a function $f(x)$ to solution $u(x)$ as follows:
@@
The Green’s function to the Poisson differential equation (DE) is:
@@
In the ML problem, the differential equation is typically unknown and we are given discrete data of forcing functions $f(x_{i})$ and solutions $u(x_{i})$. We seek to train a model to approximate a solution operator that allows us compute unknown $u$ given unseen $f$. The domain of this problem is continuous, but measuring the fields introduces a discretization: values of $u(x_{i})$ and $f(x_{i})$ corresponding to measurements of the fields on a grid $x_{i}$. In practice, this discretization arises from a sensor grid or a camera, such that each grid point is analogous to a pixel in an image. For simplicity, we only consider uniform spacing of $\Delta x$. The simplest black-box model is a linear solution operator given by
@@
where ${\bm{W}}$ is of shape $N_{\text{grid}}\times N_{\text{grid}}$ and can be trained via the usual gradient descent procedure on a mean square error loss, $\mathcal{L}=\sum\|{\bm{u}}-{\bm{W}}{\bm{f}}\|^{2}_{2}$.
A natural question arises as to what sort of matrix this model learns. We illustrate the impetus for this paper in Fig. 1. At first, we constructed a dataset using analytical solutions with cubic forcing functions, and achieved low train and test error on that space. The learned matrix (second row) qualitatively resembles the analytically derived matrix from Green’s function (first row). On further inspection, the learned matrix is quite far from the “truth”. The key insight is to take extra care on how the dataset is constructed. In general, the model will appear uninterpretable and fail to generalize to out-of-distribution (OOD) eval datasets. However, with careful construction of the dataset (bottom of Fig  1), we can actually recover both a full Green’s function discretization, ${\bm{A}}$ and a discretization of the differential operator, ${\bm{L}}={\bm{A}}^{-1}$.
To formalize, we define generalizability as achieving a low error, $\mathcal{L}()$, when sampling on a different the distribution than the training data. After minimizing $\mathcal{L}(W,\{u,f\}\sim\mathcal{F}_{train})$, we then desire:
@@
when $\mathcal{F}_{\text{test}}\neq\mathcal{F}_{\text{train}}$. The usage of $\lesssim$ denotes a small “wiggle room” tolerance heuristic. We observe that the OOD error can be orders of magnitude larger than the observed training error, $\mathcal{L}_{\text{test}}\gg 10^{8}\times\mathcal{L}_{\text{train}}$.
We design a new cross-validation procedure that trains a given ML model with a dataset on a subspace, and then evaluates the error on datasets generated on different subspaces. We construct datasets by varying the function class of solutions $f\sim\mathcal{F}(type,p)$ and the grid spacing $\Delta x$ of the discretization. We consider four different modeling approaches make the following contributions:
@@
    Underparameterized models with an inductive bias of known physical equations. We demonstrate both empirically and theoretically that the estimated $w$ parameter converges to the true parameter $k$ at a rate dependent upon the grid spacing $\Delta x$, finite difference order $q$, and which increases with the order of training function basis $p$.
@@
    Exact parametrization for linear problems. We derive a true solution ${\bm{A}}$, and theoretically and empirically demonstrate that the learned weights only converges to a subspace of the operator based on the basis of the training data. The convergence is not bound by $\Delta x$.
@@
    Overparameterized neural models. We empirically demonstrate that deep models are not guaranteed to generalize outside of the training data. Subspace generalization is also not guaranteed. It is possible to extract Green’s function operators from black-box models.
@@
    Methods designed for differential equation learning. We observe that the DeepONet and Fourier Neural Operator also do not generalize well and behave qualitatively similarly to deep linear models in our experiments.
Counterintuitively, we discover that the data requirements and generalization results for each class of model are different and even opposite in some cases.
@@ 2 Background
Applications of ML in science relevant to discovering unknown dynamics from data generally span across “white-box” to “black-box”. White-box methods, such as SINDy and symbolic regression, allow for the obtainment of closed-form, analytical formulas (Rudy et al., 2017; Udrescu & Tegmark, 2020). Black-box methods such as Neural Operators or DeepONets generally use NNs to learn dynamics (Raissi et al., 2019; Lu et al., 2019), exchanging interpretability for flexibility. Between these two approaches, techniques have been developed that respect physical constraints (Hansen et al., 2023b) such as conservation laws (Hansen et al., 2023a; Patel et al., 2022) or other fundamental identities (Trask et al., 2022). Extending even beyond these, “world models” seek to replace simulation environments (Ha & Schmidhuber, 2018) with ML models operating directly on raw state-action observations data. Massive video-action transformer models are now commonly used in deployed self-driving car systems (Goff et al., 2025) and are expanding to other domains such as surgical robotics (Schmidgall et al., 2024). These applications require high quantitative accuracy and generalization. Even a 1% error on a latent physical parameter can have a grave impact.
Interpretability of internal model mechanics has become a subject of intense interest as ML models become more capable and general purpose. The phase change in deep models between memorization to generalization has been identified as the grokking phenomenon, with similar works detailing of watching weight matrix evolution toward prescribed algorithms (Power et al., 2022; Nanda et al., 2023). In a similar vein as scientific interpretability, transformers have been shown to be able to express and converge to algorithms for linear regression Zhang et al. (2024); Liang et al. (2024).
Recent strides in numerical analysis of ML have demonstrated that Neural ODEs and residual connections can overfit to time series data and may not learn true continuous models (Queiruga et al., 2020; Ott et al., 2021; Sander et al., 2022). Furthermore, it has been proven that higher order time discretization can elicit generalization(Zhu et al., 2022; Krishnapriyan et al., 2023). For spatiotemporal models, Krishnapriyan et al. (2021) demonstrated that PINNs will fail without controlled training strategies. The usage of Green’s functions in machine learning has been examined and used to prove data requirements for elliptic equations (Gin et al., 2021; Boullé et al., 2022, 2023). Other interpretations of SciML mechanisms as computational mathematics techniques include drawing analogies between transformers and quadrature, nonlocal operators, or kernel functions (Cao, 2021; Nguyen et al., 2022; Yu et al., 2024).
@@ 3.1 Dataset construction
To explore how the origin of training data impacts the model, we construct different datasets by sampling from different function spaces, assuming no noise. Let $\mathcal{F}(type,p)$ denote a function space of a type (polynomial, FEM, cosine, sine, etc.) with $p$ terms. Given the function class, the dataset is then constructed by sampling functions and evaluating on a grid with points $x_{i}\in i\Delta x$ with $\Delta x=1/N_{\text{grid}}$ (e.g. the spacing the sensor grid). Entries in the dataset are $\{x_{i},f(x_{i}),u(x_{i})\}$ from functions sampled from $f\sim\mathcal{F}(type,p)$.
In the theoretical analysis, we consider functions constructed from basis sets. We implement four distinct function classes for experiments with variable $p$. The first class is polynomials, $f(x)=\sum_{i=1}^{p}c_{i}x^{i}$. The Cosine and Sine datasets are defined by $f(x)=\sum_{i=1}^{p}c_{i}\cos(i\pi x)$ and $f(x)=\sum_{i=1}^{p}c_{i}\sin(i\pi x)$ respectively. For these classes, functions $f$ are drawn by sampling coefficients from a distribution $c_{i}\sim U[-1,1]$, and then $u(x)$ is analytically solved. The final function class, “FEM”, is constructed from piecewise linear $f$ with sampled node values $f_{i}\sim\mathcal{N}(0,1)$. The finite element method (FEM) provides analytically exact solutions for $u(x_{i})$ at node points in 1D. In this dataset, each point value ${\bm{f}}_{i}$ is independently random. The corresponding solution $u$ is determined from this specific $f$ (Fig. 1), which is distinct from adding noise to the observations of $f$ and $u$.
3.2 Parameter learning with a numerically discretized differential equations
Now consider the case where one has identified the form of the DE, through either prior knowledge or a system identification method, but do not know its parameters. The task is to fit a discretization of the Poisson equation for the single scalar parameter $w$, whose true counter part is $k$ in (1), to satisfy the equation $-w\mathrm{d}^{2}u(x_{i})/dx^{2}\approx f(x_{i})$ over the data. The DE is approximated using a finite difference stencil with the given $\Delta x$. E.g., the three-point stencil, referred to as FD-2, is
@@ 3.2 Parameter learning with a numerically discretized differential equations
It approximates the derivative with a local truncation error proportional to the grid size at a rate $q=2$. We can estimate $w$ given the data $\left\{u,f\right\}$ by analytically solving the minimization problem of the mean square error (MSE) loss over $N$ examples
$w=\arg\min_{w}\frac{1}{N}\sum_{n=1}^{N}\sum_{i=0}^{N_{grid}}\left(\frac{w}{\Delta x^{2}}\left(u_{i-1}^{(n)}-2u_{i}^{(n)}+u_{i+1}^{(n)}\right)+f_{i}^{(n)}\right)^{2}$ (7)
where superscript $(n)$ indexes a training example. The following theorem shows that the relative error $e=|w-k|/|k|$ can be bounded under the assumption of data sampled from polynomials:
Theorem 1.
Learning the parameter $k$ using a finite difference stencil of order $q$ given polynomial training data of order $p$ on a grid of spacing $\Delta x$ results in $w=k$ when $p<q$ and an error
$\frac{|w-k|}{|k|}=\mu_{q}\Delta x^{q}+\sum_{m=q+1}^{p}\mu_{m}\Delta x^{m}\quad when\quad p\geq q$ (8)
for constants $\mu_{m}$ depending on the truncation error coefficients of the finite difference stencils.
The proof is in Appendix A.1. Even in the absence of numerical error, there is still an error between the learned parameter $w$ and true parameter $k$ based on the grid size $\Delta x$ and polynomial order of the dataset. Increasing the grid size $\Delta x$ increases the error proportional to $\Delta x^{q}$. Incrementing the polynomial order of the training data from $p-1$ to $p$ increases the error proportional to $\Delta x^{p}$.
@@ 3.3 Training Dynamics for a Linear Model
We wish to learn a matrix ${\bm{W}}$ which converges to the action of a solution operator ${\bm{A}}$. Recall the Green’s function (2) for the Poisson problem, and its discretized problem (4). It is necessary to also discretize the Green’s function relative to the discrete measurement space. The continuous function can be constructed by $f(x)=\sum_{i}{\bm{f}}_{i}\phi_{i}(x)$ where $\psi_{i}$ is a basis of the grid discretization. Basis functions $\psi_{i}(x)$ could be piecewise constant representing camera pixels, or piecewise linear, representing interpolation between point sensors. The discrete solution operator is thus,
[…]
We let ${\bm{A}}_{ij}$ be the matrix in the brackets such that ${\bm{u}}={\bm{A}}{\bm{f}}$. The MSE loss over the dataset for this model is
$\mathcal{L}=\frac{1}{2N}\sum_{n=1}^{N}\left\|{\bm{W}}{\bm{f}}^{(n)}-{\bm{u}}^{(n)}\right\|^{2}_{2}$ (10)
We consider the case where our data is only sampled uniformly over subspaces of $\mathbb{R}^{n}$, such that ${\bm{B}}$ is the matrix of rank $p+1$ of coefficients satisfying $f(x_{i})={\bm{f}}_{i}={\bm{B}}{\bm{c}}$. For example, a polynomial function space can use the Vandermonde matrix ${\bm{B}}_{ij}=[1,x_{i},x_{i}^{2}...x_{i}^{p}]$. By Eq. (9), there exists a matrix ${\bm{A}}$ such that all training and test data satisfies ${\bm{u}}={\bm{A}}{\bm{f}}$. The following result shows that out of distribution generalizations cannot be expected:
Theorem 2.
Applying gradient descent a training dataset $\{{\bm{u}}^{(n)},{\bm{f}}^{(n)}\}$ sampled from a function space $\mathcal{F}_{\text{train}}$, the model weights ${\bm{W}}$ will converge to projection of the true operator ${\bm{A}}$ onto the subspace of $\mathcal{F}_{\text{train}}$. Let the forcing functions be sampled by ${\bm{f}}^{(n)}={\bm{B}}{\bm{c}}^{(n)}$ where each component of ${\bm{c}}$ is sampled independently with $\mathbb{E}[{\bm{c}}_{i}]=0$. Optimizing the loss (10) using gradient descent with an initial condition ${\bm{W}}^{0}$ will converge to
@@
The proof is in Appendix A.3. The learned matrix will converge to the projection of the true operator ${\bm{A}}$ onto the subspace of data $F$, ${\bm{A}}{\bm{U}}{\bm{U}}^{T}$, plus an error/noise that is orthogonal to the subspace of the data relative to the initial weights ${\bm{I}}-{\bm{U}}{\bm{U}}^{T}$. The latter “noisy” aspect could be mitigated by regularization such as weight decay, converging to ${\bm{W}}={\bm{A}}{\bm{U}}{\bm{U}}^{T}$ but not ${\bm{A}}$ exactly.
The above result is irrespective of the amount of data and the discretization, and as such, is actually a quite pessimistic result. In particular, ${\bm{A}}$ is only learned if and only if the dimension of subspace of ${\bm{f}}$ is of equal rank with ${\bm{A}}$. The finer the discretization, the greater the “quality” of the data needed (e.g. dimension of the subspace); infinite quantity of data is not sufficient.

### v2

@@ Abstract
While there are many applications of machine learning (ML) to scientific problems that look promising, the eye test can be misleading compared to quantitative measures. Using numerical analysis techniques, we rigorously quantify the accuracy, convergence rates, and generalization bounds of certain ML models applied to linear differential equations (DEs) for parameter discovery or solution finding. Beyond the quantity and discretization of data, we identify that the function space of the data is critical to the generalization of the model. A similar lack of generalization is empirically demonstrated for commonly used models, including physics-specific techniques. Counterintuitively, we find that different classes of models can exhibit opposing generalization behaviors. Based on our theoretical analysis, we also introduce a new mechanistic interpretability lens on scientific models whereby Green’s function representations can be extracted from the weights of black-box models. Our results inform a new cross-validation technique for measuring generalization in physical systems, which can serve as a benchmark.
@@ 1 Introduction
The development of robust and rigorous a priori estimates for numerical methods is a major victory of the past half-century. These estimates give geometric flexibility and convergence guarantees, laying the groundwork for complex physics simulations across a wide range of applications. The basis of the theory begins with the humble 1D Poisson equation, studied by virtually all numerical analysts and serving as the bedrock for computational physics and engineering.
The canonical Poisson equation on the unit interval with homogeneous Dirichlet boundary conditions is given by
$k\frac{d^{2}u}{dx^{2}}=f(x),\qquad u(0)=u(1)=0.$ (1)
where $u(x),f(x),k$ are the solution, forcing function, and (constant) material coefficient, respectively. The solution to Eq. (1) can be obtained using the Green’s function which maps a function $f(x)$ to the solution $u(x)$ as follows:
@@
where
@@
In the ML context, the governing DE is typically unknown, and we are given discrete evaluations of forcing functions $f(x_{i})$ and solutions $u(x_{i})$ with $x_{i}\in(0,1)$ corresponding to a sensor grid or pixels from a camera. We then seek to train a model approximating the solution operator that can compute an unknown $u$ given an unseen $f$. While the DE is continuous, the act of measurement discretizes all objects. For simplicity, we only consider uniformly spaced points $x_{i}=i\Delta x$ for $i=0,\ldots,N_{\text{grid}}$.
The simplest black-box model is a linear operator given by
@@
where ${\bm{W}}\in\mathbb{R}^{N_{\text{grid}}\times N_{\text{grid}}}$ and is trained via the usual gradient descent on mean squared error (MSE) loss, $\mathcal{L}=\sum\|{\bm{u}}-{\bm{W}}{\bm{f}}\|^{2}_{2}$. A natural question is what matrix this model learns; see Fig. 1. First, we constructed a dataset using analytical solutions with cubic forcing functions, and achieved low train and test error on that space. The learned matrix (second row) qualitatively resembles the analytically derived matrix from Green’s function (first row), but is quite far from the “truth” when we look at the inverse ${\bm{W}}^{-1}_{p=3}$. In general, the model will appear uninterpretable and fail to generalize to out-of-distribution (OOD) datasets. However, with careful construction of the dataset (bottom of Fig  1), we can recover both a full Green’s function discretization, $\bm{A}$} and a discretization of the differential operator, ${\bm{L}}={\bm{A}}^{-1}$.
Figure 1: Expectations when learning a black box linear model, ${\bm{u}}={\bm{W}}{\bm{f}}$. Despite some visual similarity to the Green’s function operator (top), naive approaches are not guaranteed to converge to the true operator. The issue lies in the sampling procedure of the training data. The naive approach of using polynomial analytical solutions (illustrated as cubic polynomials) can achieve machine precision MSE loss, but even with infinitely many examples, will not converge to the general solution (center). Altering the construction of the training data, i.e. by using random piecewise linear functions (bottom) can recover the true operator. This allows for the extraction of a familiar discrete differential equation operator (right).
To formalize, we define generalizability as achieving a low error when sampling from a different distribution than the training data. After minimizing $\mathcal{L}(W,\{u,f\}\sim\mathcal{F}_{\text{train}})$, we then desire that the model satisfy:
@@
when $\mathcal{F}_{\text{test}}\neq\mathcal{F}_{\text{train}}$. The symbol $\lesssim$ denotes a small “wiggle room” tolerance heuristic. We empirically observe that the OOD error can be orders of magnitude larger than the observed training error, $\mathcal{L}_{\text{test}}\gg 10^{8}\times\mathcal{L}_{\text{train}}$, which makes the OOD transition apparent.
We also design a new cross-validation procedure that trains a given ML model with a dataset on a subspace, and then evaluates the error on datasets generated on different subspaces. The datasets are varied not only by the grid spacing $\Delta x$, but also by the function class $f\sim\mathcal{F}(type,p)$. We consider five different approaches and make present following contributions:
@@
    Underparameterized models with an inductive bias of known physical equations — We demonstrate both empirically and theoretically that the estimated $w$ parameter converges to the true parameter $k$ at a rate dependent upon the grid spacing $\Delta x$, finite difference order $q$, and which increases with the order of training function basis $p$.
@@
    Exact parametrization for linear problems — We derive a true solution $\bm{A}$}, and theoretically and empirically demonstrate that the learned weights only converge to a subspace of the operator based on the basis of the training data. The grid size $\Delta x$ does not change the error, but changes the scaling of learned parameters.
@@
    Overparameterized neural models — We empirically demonstrate that deep models are not guaranteed to generalize outside of the training data. Subspace generalization is also not guaranteed. It is possible to extract Green’s function operators from black-box models.
@@
    Methods designed for DE learning — We observe that the DeepONet and Fourier Neural Operator also do not always generalize, similarly to black-box models.
  5. 5.
    Physics informed losses — We demonstrate that PINNs and Physics Informed DeepONets, which incorporate prior DE knowledge, exhibit similar failure modes.
Surprisingly, we find that the data requirements and generalization behavior for each model class can vary dramatically, and in some cases, are even contrary to one another.
@@ 2 Background
Applications of ML in science span across “white-box” to “black-box”. White-box methods, such as SINDy and symbolic regression, allow for the obtainment of closed-form, analytical formulas (Rudy et al., 2017; Udrescu and Tegmark, 2020). Black-box methods such as Neural Operators or DeepONets generally use NNs to learn dynamics (Raissi et al., 2019; Lu et al., 2019), exchanging interpretability for flexibility. Between these two approaches, techniques have been developed that respect physical constraints such as conservation laws (Hansen et al., 2023; Patel et al., 2022; Trask et al., 2022). However, the adoption of these methods in high-consequence scenarios is lacking due to not understanding the exact failure modes.
To this end, recent studies have demonstrated that Neural ODEs and residual connections can overfit to time series data and may not learn true continuous models (Queiruga et al., 2020; Ott et al., 2021; Sander et al., 2022), whereas higher order time discretization can elicit generalization (Zhu et al., 2022; Krishnapriyan et al., 2023). For spatiotemporal models, Krishnapriyan et al. Krishnapriyan et al. (2021) demonstrated that PINNs will fail without controlled training strategies, whereas Sakarvadia et al. Sakarvadia et al. (2025) showed that PINOs fail to generalize across spatial resolutions. The usage of Green’s functions in machine learning has been examined and used to prove data requirements for elliptic equations (Gin et al., 2021; Boullé et al., 2022, 2023). Direct inspection of of weight matrices is an active area of research in mechanistic interpretability, where known algorithms can be found in the weights of ML models (Nanda et al., 2023). Interpretations of SciML mechanisms include drawing analogies between transformers and quadrature, nonlocal operators, or kernel functions (Cao, 2021; Nguyen et al., 2022; Yu et al., 2024).
@@ 3.1 Dataset construction
To explore how the functional source of training data impacts the model, we construct different datasets by sampling from different function spaces. Let $\mathcal{F}(type,p)$ denote a function space of a type (e.g., polynomial, FEM, cosine, sine, etc.) with $p$ terms. Given the function class, the dataset is constructed by sampling functions evaluated on a grid with points $x_{i}\in i\Delta x$ with $\Delta x=1/N_{\text{grid}}$ (e.g. the spacing the sensor grid). Entries in the dataset are $\{x_{i},f(x_{i}),u(x_{i})\}$ from functions sampled from $f\sim\mathcal{F}(type,p)$. In experiments, we consider function types of polynomial, sine $\sin(k\pi x)$, cosine $\cos(k\pi x)$, and piecewise linear.
Figure 2: Comparison of numerical experiments to analytically derived ML parameters from Theorems 3.1 and 3.2 trained on polynomial spaces with $N_{grid}=22$. (Left) The observed error for the linear matrix matches predictions for lower order pp. (Middle and right) For the finite difference model, the analytical assumptions overestimate the error, but the trend matches with observations.
Figure 3: Learning a parameter using a PINN inverse problem. Line is the mean of five runs, and the shaded region covers the min-max spread.
3.2 Parameter learning with a discretized DE
Consider the case where the form of the DE is known, through either prior knowledge or a system identification method, but its physical parameters are unknown. We fit the discretized Poisson equation for a parameter $w$ (whose true value is $k$ from Eq. 1) such that $w\frac{d^{2}u(x_{i})}{dx^{2}}\approx f(x_{i})$ over the data. The DE is approximated using a standard finite difference stencil, such as the three-point stencil (FD-2),
@@ 3.2 Parameter learning with a discretized DE
We can estimate $w$ given the data $\left\{u,f\right\}$ by analytically minimizing the MSE loss over $N$ examples
$\frac{1}{N}\sum_{n=0}^{N}\sum_{i=1}^{N_{grid}-1}\left(\frac{w}{\Delta x^{2}}\left(u_{i-1}^{(n)}-2u_{i}^{(n)}+u_{i+1}^{(n)}\right)+f_{i}^{(n)}\right)^{2}$ (7)
where superscript $(n)$ indexes training examples. The following theorem shows that the relative error e=|wk|/|k|e=|w-k|/|k| can be bounded for data sampled from polynomials:
Theorem 3.1.
Learning the parameter $k$ using a finite difference stencil of order $q$ given polynomial training data of degree $p$ on a grid of spacing $\Delta x$ results in $w=k$ when $p<q$, and an error
$\frac{|w-k|}{|k|}=\mu_{q}\Delta x^{q}+\sum_{m=q+1}^{p}\mu_{m}\Delta x^{m}\approx\mu_{q}\Delta x^{q}\,\$(8)
when $p\geq q$, for constants $\mu_{m}$ depending on the truncation error coefficients of the finite difference stencil.
The proof is in App. A.1. Even without numerical error or noise, the learned parameter $w$ has an error depending on the stencil order. Counterintuitively, increasing the polynomial degree of the dataset is counterproductive.
@@ 3.3 Training Dynamics for a Linear Model
Suppose now, we wish to learn a matrix $\bm{W}$} which converges to the action of a solution operator $\bm{A}$}. To determine the expected $\bm{A}$}, the Green’s function needs to be discretized relative to the discrete measurement space. The continuous forcing function can be constructed by $f(x)=\sum_{i}{\bm{f}}_{i}\psi_{i}(x)$ where $\psi_{i}$ is a basis of the grid discretization. The discrete solution operator is thus,
${\bm{u}}_{i}=\sum_{j}\left[\int_{0}^{1}G(x_{i},s)\psi_{j}(s)\mathrm{d}s\right]{\bm{f}}_{j}$ (9)
with the matrix in brackets defines $\bm{A}$}. The MSE loss over the dataset for this model is
$\mathcal{L}=\frac{1}{N}\sum_{n=1}^{N}\left\|{\bm{W}}{\bm{f}}^{(n)}-{\bm{u}}^{(n)}\right\|^{2}_{2}$ (10)
More generally, suppose the data is sampled uniformly over subspaces of $\mathbb{R}^{n}$ such that $\bm{B}$} is the matrix of rank $p+1$ of coefficients satisfying $f(x_{i})={\bm{f}}_{i}={\bm{B}}{\bm{c}}$. For example, a polynomial function space can use the Vandermonde matrix ${\bm{B}}_{ij}=[1,x_{i},x_{i}^{2}...x_{i}^{p}]$. By (9), there exists a matrix $\bm{A}$} such that all training and test data satisfies ${\bm{u}}={\bm{A}}{\bm{f}}$. The following result shows that out of distribution generalizations cannot be expected:
Theorem 3.2.
Applying gradient descent on training dataset $\{{\bm{u}}^{(n)},{\bm{f}}^{(n)}\}$ sampled from a function space $\mathcal{F}_{\text{train}}$, the model weights $\bm{W}$} will converge to projection of the true operator $\bm{A}$} onto the subspace of $\mathcal{F}_{\text{train}}$. Let the forcing functions be sampled by ${\bm{f}}^{(n)}={\bm{B}}{\bm{c}}^{(n)}$ where each component of $\bm{c}$} is sampled independently with $\mathbb{E}[{\bm{c}}_{i}]=0$. Optimizing the loss (10) using gradient descent with an initial condition ${\bm{W}}^{0}$ will converge to
@@
The proof is in App. A.3. The learned matrix will converge to the projection of the true operator $\bm{A}$} onto the subspace of data $F$, ${\bm{A}}{\bm{U}}{\bm{U}}^{T}$, plus an initial error/noise that is orthogonal to the subspace of the data. This result is irrespective of the amount of data and the discretization, and as such, is actually a quite pessimistic result. In particular, $\bm{A}$} is only learned iff the dimension of subspace of $\bm{f}$} is of equal rank with $\bm{A}$}. The finer the discretization, the greater the “quality” of the data needed (e.g. dimension of the subspace).